OpenStudy (elusive):

In a 30-60-90 right triangle, the length of the hypotenuse is ________ times the length of the short leg, and the length of the long leg is _______ times the length of the short leg

OpenStudy (thephysicsman):

let's say if the hypothenuse was 2 and the opposite side was 1 then we can easily find the value of a the other leg.

OpenStudy (thephysicsman):

\[a^{2}+b^{2} =c^{2}\] \[\sqrt{ c^{2}-a^{2}} = b => \sqrt{4-1} => \sqrt{3}\] \[\sin(30^{o}) = \frac{ opposite }{ hypothenuse } = \frac{ 1 }{ 2 } = \frac{ a }{ c }\]  so just knowing this you can take a look at what we have here.